OpenStudy (anonymous):

A bag contains 3 coins, 1 is baised with both side heads and other two are normal. Assume 'A' to be baised, ['B' and 'C'] non baised. A coin is chosen at random from the bag and tossed 4 times. If heads turn up each time, what is the prob that it is a 2 headed coin.

OpenStudy (anonymous):

put \(Y\) as the event the coin is two headed, \(X\) as the event you get 4 heads, and you need to compute \[P(Y|X)=\frac{P(Y\cap X)}{P(X)}\] each coin is chosen with probability \(\frac{1}{3}\)

OpenStudy (anonymous):

I know i will have to use Baye's theorem in this... prob that A, B or C turns up is P(A)=P(B)=P(C)=1/3 Prob that heads turns from A, B or C will be P( H|A) =1

OpenStudy (anonymous):

the real work is computing \(P(X)\)

OpenStudy (anonymous):

I want to know why Prob of P( H|B) =\[(\frac{ 1 }{ 2 }) ^{4}\]

OpenStudy (anonymous):

if it is an unbiased coin, then the probability you get 4 heads is \(\frac{1}{16}\) so the probability you get 4 heads is \[\frac{1}{3}\times \frac{1}{16}+\frac{1}{3}\times \frac{1}{16}+\frac{1}{3}\times 1\]

OpenStudy (anonymous):

the power 4 is there coz of coin being tossed 4 times??

OpenStudy (anonymous):

i.e. coin B is chosen, AND get 4 heads OR coin C is chosen AND get 4 heads OR coin A is chosen AND get 4 heads

OpenStudy (anonymous):

yes, the probability you get 4 heads on an unbiased coin is \(\left(\frac{1}{2}\right)^4=\frac{1}{16}\)

OpenStudy (anonymous):

numerator of your fraction is \(\frac{1}{3}\) and denominator is \(\frac{1}{3}\times \frac{1}{16}+\frac{1}{3}\times \frac{1}{16}+\frac{1}{3}\times 1\) whatever that is

OpenStudy (anonymous):

OK can u suggest me a good book where I can find solution or solved examples for probability??

OpenStudy (anonymous):

depends on what level you want Feller used to be standard text there is a good book by Galombos and one on discrete probability that i like by Gordon but I would look for free on line resources

OpenStudy (anonymous):

I am 2nd year college student.

OpenStudy (anonymous):

I am reffering Meyer

OpenStudy (anonymous):

look on line

OpenStudy (anonymous):

online as in Yahoo answers??

OpenStudy (anonymous):

i like google there are no doubt free on line texts

OpenStudy (anonymous):

thank you